OpenStudy (anonymous):

HI! I do get difficulties on how to perform these partial derivatives qns. EG,If z =f(x,y) is a differentiable f of x and y. where x=rcosA and y=rsinA. my problem is that how to get dz/dr while in the problem there is no more elaborations of z rather than z=f(x,y)?

OpenStudy (anonymous):

F(x,y,z)=f(x,y)-z

OpenStudy (anonymous):

well remember chain rule for partial differentiation? suppose if we have z=f(x,y), and x=f(r,A) and y=f(r,A), so dz/dr is given by dz/dr=(dz/dy)(dy/dr)+(dz/dx)(dx/dr)=(dz/dy)(sinA)+(dz/dx)(cosA)

OpenStudy (anonymous):

if there's no further evaluation of z, I'm afraid this will only be dz/dr

OpenStudy (anonymous):

thanks very much anonymoustwo44, your advice worked, I am real enjoying the questions now.